# Construction of Solutions with Extraordinary Gradient Amplification and Localization for Schrödinger Equations

### v1

@@ Abstract.
2020 Mathematics Subject Classification:  35Q55,35B65,35P25,35B40,47A75,35B30,81Q10
@@
E-mail: o@jlu.edu.cn, hadiao@gmail.com
@@ 1. Introduction
1.1. Mathematical setup and summary of major results
@@
Initially we focus on the mathematical models for both linear and nonlinear Schrödinger equations in two and three spatial dimensions in this paper, where the mechanism for these mathematical models will be elaborated in more detail later. Let the unknown function $u:\Omega\times[0,T]\rightarrow\mathbb{C}$ satisfy the homogeneous initial-boundary value problem for a linear Schrödinger equation:
@@
$\begin{cases}\mathrm{i}\partial_{t}u+L_{1}u=0&\text{in }\Omega\times(0,T],\\ u(\mathbf{x},0)=\phi(\mathbf{x})&\text{in }\Omega,\\ u(\mathbf{x},t)=\psi(\mathbf{x},t)&\text{on }\partial\Omega\times[0,T],\end{cases}$ (1.1)
@@
and let $U:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{C}$ satisfy the Cauchy problem for a nonlinear Schrödinger equation:
@@
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}U+L_{2}U+N(U)=0&&\text{in }\mathbb{R}^{d}\times(0,T],\\ &U(\mathbf{x},0)=\Phi(\mathbf{x})&&\text{in }\mathbb{R}^{d},\end{aligned}\right$. (1.2)
where the following conditions hold:
@@
    $\mathrm{i}$ denotes $\sqrt{-1}$, $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{d}$, $d=2,3$, and $0<T<\infty$.
  2. (2)
    $L_{1}=\nabla\cdot(\mathbf{A}_{1}\nabla)$, $L_{2}=\nabla\cdot(\mathbf{A}_{2}\nabla)+c$.
  3. (3)
    The complex-valued coefficient $\mathbf{A}_{1}=(a_{1}^{ij}(\mathbf{x}))_{i,j=1}^{d}$ satisfies:
    1. (a)
      $\overline{\mathbf{A}}_{1}^{\top}\!(\mathbf{x})=\mathbf{A}_{1}(\mathbf{x}),~\forall\mathbf{x}\in\Omega$;
    2. (b)
      $\overline{\xi}^{\top}\mathbf{A}_{1}(\mathbf{x})\xi\geq\theta\lvert\xi\rvert^{2}$ for a.e. $\mathbf{x}\in\Omega$, $\forall\xi\in\mathbb{C}^{d}$, with constant $\theta>0$;
    3. (c)
      $\mathbf{A}_{1}\in W^{2,\infty}(\Omega)\cap C^{1,1}(\overline{\Omega})$ ;
    where $\overline{\cdot}$ denotes the complex conjugate and $\top$ denotes the transpose.
  4. (4)
    $D$ is a bounded Lipschitz domain such that $D\Subset\Omega$, $\D¯\Omega\backslash\overline${D} is connected and $\mathbf{A}_{1}$ has compact support in $D$, i.e.,
    $\operatorname{supp}(\mathbf{A}_{1}-\mathbf{I})\subset D.$
  5. (5)
    The complex-valued coefficients $\mathbf{A}_{2}(\mathbf{x},t)=\big(a_{2}^{ij}(\mathbf{x},t)\big)_{i,j=1}^{d},c(\mathbf{x},t)$ satisfy the following conditions:
    1. (a)
      The matrix $\mathbf{A}_{2}(\mathbf{x},t)$ is given by
      $\mathbf{A}_{2}(\mathbf{x},t)=\begin{cases}\mathbf{B}(t),&(\mathbf{x},t)\in\overline{D}\times[0,T],\\[4.0pt] \mathbf{I},&(\mathbf{x},t)\in(\mathbb{R}^{d}\setminus\overline{D})\times[0,T],\end{cases}$
      where $\mathbf{I}$ is the $d$-dimensional identity matrix and $\mathbf{B}(t)=\big(b^{ij}(t)\big)_{i,j=1}^{d}$ is a Hermitian matrix and $\mathbf{B}\in L^{\infty}[0,T]$. Moreover, it satisfies the uniform ellipticity condition
    2. (b)
      The function $c\in L^{\infty}\bigl(0,T;W^{3,\infty}(\mathbb{R}^{d})\bigr),$ and its spatial support is contained in $D$ for all $t\in[0,T]$, i.e.,
      $\operatorname{supp}c(\cdot,t)\subset D,\qquad\forall t\in[0,T].$
  6. (6)
    The nonlinearity $N:\mathbb{C}\to\mathbb{C}$ is a complex polynomial given by
    $N(U)=\sum_{k=2}^{l_{0}}\alpha_{k}U^{k},$ (1.3)
    where $l_{0}\geq 2$ is a fixed integer. The coefficients $\alpha_{k}$ satisfy::
    with compact support in $D$, i.e.,
  7. (7)
    The initial data $\phi\in C^{\infty}{(\Omega)}$, $\Phi\in C^{\infty}(\mathbb{R}^{d})$, and boundary data $\psi\in C^{\infty}([0,T];C(\partial\Omega))$ will be specified in Section 3.
In this paper, we identify and construct a mechanism that produces highly localized gradient amplification of solutions to both the linear and nonlinear Schrödinger-type equations in two and three spatial dimensions. The phenomenon is described by the following two features:
  1. (i)
    For any prescribed set of $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, where $D$ is the support of the anisotropic coefficients $\mathbf{A}_{1}$ in (1.1) or $\mathbf{A}_{2}$ in (1.2), the zero-order term $c$ in (1.2), and the nonlinear term $N$ in (1.2), we rigorously prove—by selecting suitably tuned smooth initial–boundary data for (1.1) or smooth initial data for (1.2)—that there exist localized gradient amplification regions in neighborhoods of each $\mathbf{x}_{i}$ located outside $D$ ($i=1,\ldots,n$). In these regions, the $C^{0}$-norms of the spatial gradients of the solutions, namely $\nabla u(t)$ to (1.1) and $\nabla U(t)$ to (1.2), exceed any prescribed large constant $\mathcal{M}$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\alpha}$-norm of the solution in these regions and the $C^{1,\alpha}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$. These properties are established in Theorems 1.1 and 1.2, where the time horizon $T\in\mathbb{R}_{+}$ is arbitrary.
  2. (ii)
We are now in a position to state our main theorems. For initial-boundary value problem of a linear Schrödinger-type equation (1.1), we have:
Theorem 1.1.
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial and boundary inputs ((as specified in (3.21))), there exists a small parameter
$r_{0}=\min\left\{\,\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|,\ \frac{1}{3}\operatorname{dist}(D,\partial\Omega)\,\right\}$
such that problem (1.1) admits a unique solution $u$ satisfying ::
$u\in L^{\infty}\big([0,T];H^{3}_{\mathrm{loc}}(\Omega)\big),$
with the following properties ::
  1. (1)
    Interior improved Hölder regularity :: for any bounded Lipschitz domain $\Omega^{\prime}\Subset\Omega$,
@@
    Vanishing measure of the localized gradient amplification set :: for almost every $t\in[0,T]$,
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla u(\mathbf{x},t)|>\mathcal{M}\right\}$
[…]
    as $\mathcal{M}\to\infty$.
Similarly, for Cauchy problem of a nonlinear Schrödinger-type equation (1.2), we have:
@@
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial inputs ((as specified in (3.22))), there exists a small parameter
$r_{0}=\min\left\{\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|\right\}$
such that problem (1.2) admits a unique solution $U$ satisfying:
@@
with the following properties::
@@
    Hölder regularity: For any bounded Lipschitz domain $\Omega^{\prime}\subset\mathbb{R}^{d}$, U L^([0,T]; C^1,12(Ω’)).
@@
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla U(\mathbf{x},t)|>\mathcal{M}\right\}$
[…]
    as $\mathcal{M}\to\infty$.
The construction mechanism developed in this work for producing customized solutions to Schrödinger equations with localized gradient amplification exhibits the following key features.
First, it applies to a broad class of linear and nonlinear Schrödinger-type equations in two and three spatial dimensions. Inside the region $D$, the coefficients may be quite general (e.g., smooth anisotropic coefficients or nonlinearities given by arbitrary polynomials), while outside $D$ the equation coincides with the standard Schrödinger equation.
Second, for any prescribed time horizon $T\in\mathbb{R}_{+}$, any arbitrarily large amplification threshold $\mathcal{M}>0$, and any finite collection of distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, one can construct suitably designed smooth initial and/or boundary data such that, for both the linear problem (1.1) and the nonlinear problem (1.2), the magnitude of the spatial gradient of the corresponding solution exceeds $\mathcal{M}$ in neighborhoods of each prescribed point located outside $D$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\alpha}$-norm of the solution in these neighborhoods and the $C^{1,\alpha}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$. The time interval $T$ can be chosen arbitrarily.
Third, the regions where the gradient amplification occurs are contained in the balls $B_{3r_{0}}(\mathbf{x}_{i})$. As the aforementioned bound $\mathcal{M}\to\infty$, the Lebesgue measure of these regions tends to zero, in agreement with the well-posedness of the underlying problems. This illustrates a clear trade-off: achieving larger gradient magnitude requires stronger spatial concentration, whereas milder amplification allows the region to occupy a larger set. Moreover, when the points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ lie close to one another, the corresponding amplification regions must remain sufficiently confined to avoid interference.
To the best of our knowledge, this work provides the first systematic and customized construction of solutions exhibiting localized gradient amplification for Schrödinger-type equations in both linear and nonlinear regimes. A more detailed discussion of the underlying ideas and implications is given in Subsection 1.3. We conclude this subsection with several remarks concerning the methodology employed in the proofs and the relationship between the assumptions and conclusions of Theorems 1.1 and 1.2.
@@
The localized gradient amplification mechanism developed in this work for Schrödinger equations is rooted in spectral analysis, in particular the use of transmission eigenfunctions associated with the interior transmission problem [5] and their approximation by Herglotz wave functions. By exploiting these tools, we construct a carefully designed auxiliary function $u_{0}$ (see Theorem 3.1), which serves as the fundamental building block for prescribing the initial and boundary inputs in both the linear and nonlinear Schrödinger equations. Further details of this construction are provided in Section 3.
@@
Both the solution $u(t)$ to the linear problem (1.1) and the solution $U(t)$ to the nonlinear problem (1.2) exhibit local $H^{3}$-regularity for almost every $t\in[0,T]$. In the linear case, this regularity follows from the interior elliptic regularity theorem applied to $u(t)$. In the nonlinear setting, the regularity of $U(t)$ is ensured by the smoothness assumptions on the coefficients together with a Banach fixed point argument. Consequently, in both cases, the Sobolev embedding theorem implies that the solutions possess $C^{1,\frac{1}{2}}$-regularity in neighborhoods of the localized gradient amplification points for almost every $t\in[0,T]$, which guarantees that the spatial gradients are well defined in the classical sense. Moreover, if additional time-regularity assumptions are imposed on the coefficients, the above regularity results can be strengthened to hold for all $t\in[0,T]$.
Remark 1.5.
In this work, the localized gradient amplification regions are described using balls in 3D and disks in 2D. It is plausible that other types of sets could be employed to characterize such highly localized gradient regions, although doing so would require more refined technical arguments.
Remark 1.6.
@@
For both the linear problem (1.1) and the nonlinear problem (1.2), the time $T\in\mathbb{R}_{+}$ of the interval $[0,T]$ can be chosen arbitrarily large while still guaranteeing the localized gradient amplification of the solutions. In the linear case, this follows directly from the classical well-posedness theory for linear Schrödinger equations. In the nonlinear case, however, such long-time existence requires a more refined analysis. By carefully tuning the initial data and combining spectral analysis with an application of the Banach fixed-point theorem, we are able to extend the solution to (1.2) over any prescribed time interval $T$. For further details, see Remark 5.3 and the proof of Theorem 1.2.
1.2. Existing studies and motivation
We begin by recalling several classical results on singularity formation in the nonlinear Schrödinger equation
$iu_{t}=\Delta u+\mu|u|^{p-1}u,\qquad x\in\mathbb{R}^{d},\ t>0,\qquad u(x,0)=u_{0}(x).$
Blow-up refers to the existence of a finite time $T>0$ such that a suitable norm of the solution $u(t)$ becomes unbounded as $t\to T$. In the $L^{2}$-supercritical case $p>1+4/d$, Glassey [11] showed that for the focusing equation $(\mu=-1)$, smooth initial data with sufficient decay at infinity and satisfying certain energy and momentum conditions lead to finite-time blow-up. In the $L^{2}$-subcritical case $1<p<1+4/d$, global well-posedness holds for the defocusing equation $(\mu=1)$ for all finite-energy data, whereas in the focusing case $(\mu=-1)$ global existence is guaranteed below specific mass-energy thresholds [10]. In the $L^{2}$-critical case $p=1+4/d$, Merle [16] demonstrated that solutions can blow up at any prescribed finite set of points, with the blow-up profile concentrating around these locations while vanishing elsewhere. Further refinements of multi-point blow-up, including log–log type concentration, are given in [8]. These results highlight the remarkable flexibility of singularity formation in the critical regime, where spatial localization can be prescribed precisely. For general background on the nonlinear Schrödinger equation, see [20, 19, 4].
In the present work, we establish an analogous, yet fundamentally different, phenomenon for certain linear and nonlinear Schrödinger equations, where solutions remain local $H^{3}$-regular over the entire interval $[0,T]$ but nevertheless develop sharply localized regions of high spatial gradient at arbitrarily selected points. This controlled formation of high-gradient regions draws conceptual inspiration from localization phenomena extensively studied in both mathematical and physical Schrödinger theory. In systems with random or quasi-periodic potentials, Anderson localization [2, 6] demonstrates that eigenfunctions may concentrate strongly around specific spatial locations even in the absence of nonlinear effects. Such behavior plays a fundamental role in quantum scattering, wave propagation in disordered media, and the spectral analysis of Schrödinger operators. More generally, classical quantum scattering theory shows that potential landscapes, boundary conditions, and geometric trapping can generate highly localized wave functions, resonances, or tunneling effects [21, 24, 17], which govern transport and energy distribution in quantum systems.
From a physical perspective, the localization mechanism uncovered in this work is also consistent with the qualitative content of the Heisenberg uncertainty principle [12]. The uncertainty principle asserts a fundamental trade-off between spatial localization and frequency (or momentum) concentration, typically formulated in a global $L^{2}$-based phase-space framework. In our setting, although no probabilistic or Fourier-analytic uncertainty inequality is invoked, an analogous trade-off emerges at the level of deterministic PDE dynamics: as the spatial gradients of the solution become arbitrarily large near prescribed points, the regions in which this amplification occurs necessarily shrink in measure. Thus, extreme spatial localization is accompanied by increasingly large gradients, reflecting a complementary manifestation of uncertainty-type behavior within Schrödinger-type evolutions. Our results therefore provide a rigorous and deterministic illustration of how localization and high-frequency effects can coexist in regular solutions, bridging ideas from PDE analysis, quantum scattering, and foundational principles of quantum mechanics.
1.3. Main contributions and technical discussions
Our work constructs novel solutions for a broad class of Schrödinger-type equations defined by their behavior inside and outside a bounded domain $D$. Specifically, we consider Schrödinger equations that reduce to the standard Schrödinger equation outside $D$, while inside $D$ they are governed by variable-coefficient operators and may include arbitrary polynomial nonlinearities. These solutions with localized gradient amplification behavior exhibit the following precise characteristics:
  1. (i)
    For any prescribed $\mathcal{M}>0$ and any prescribed number $n$ of distinct points $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ on $\partial D$, we construct inputs (specifically, initial and/or boundary inputs) that yield solutions $u$ and $U$ to problems (1.1) and (1.2), respectively. In the specified regions near these points, the $L^{\infty}$-norms of $\nabla u$ and $\nabla U$ exhibit localized gradient amplification bounded below by $\mathcal{M}$ for almost every time $t$ in the interval $[0,T]$, where $T$ can be chosen arbitrarily in advance. Moreover, the amplification is quantitatively contrasted with the behavior inside $D$: the ratio between the local $C^{1,1/2}$-norm of the solution in each amplification region outside $D$ and the $C^{1,1/2}$-norm of the solution restricted to $\overline{D}$ is bounded from below by $\mathcal{M}/2$ for almost every $t\in[0,T]$. This demonstrates that the observed amplification is not merely absolute, but also relative to the interior regularity of the solution. For the nonlinear problem (1.2), the solution exists for an arbitrarily large time interval. This long-time existence is achieved through a combination of spectral analysis, application of the Banach fixed-point theorem, and careful tuning of the initial data (see Remark 5.3 and the proof of Theorem 1.2).
  2. (ii)
    Due to the well-posedness of problems (1.1) and (1.2), the measure of the localized gradient amplification regions—where the gradient of solutions is bounded below by $\mathcal{M}$—vanishes for almost every $t\in[0,T]$ as $\mathcal{M}\to\infty$. We provide detailed characterizations of the measure of these regions in Theorems 1.1 and 1.2. Stronger localized gradient amplification corresponds to smaller spatial support of these regions. This quantitative decay property of the measure with respect to $\mathcal{M}$ contrasts with classical blowup, where the singular set typically has positive measure at blowup time. Importantly, under the specified smooth and finely tuned inputs—even with the weaker regularity coefficients specified in Items 3 and 5 of (1.1) or (1.2) whose supports lie in a bounded Lipschitz domain—the solutions $u$ and $U$ retain local spatial $H^{3}$ regularity for almost every $t\in[0,T]$. Moreover, within the localized gradient amplification regions, the solutions maintain local spatial $C^{1,1/2}$ regularity for almost every $t\in[0,T]$.
This localized gradient amplification behavior is realized through a careful combination of spectral theory for transmission eigenfunctions of the interior transmission problem [5] and Herglotz wave function approximations. By analyzing the spectral properties of these eigenfunctions, we identify key structures that allow for the selective amplification of gradients near pre-specified points. Herglotz approximations are then employed to explicitly construct smooth initial and/or boundary inputs that excite these structures, effectively transferring the amplification to the solutions of both the linear problem (1.1) and the nonlinear problem (1.2). The resulting inputs, detailed in (3.21) and (3.22), guarantee that the gradient of the solution attains the prescribed bound $\mathcal{M}$ in the targeted regions, while maintaining regularity.

### v2

@@ Abstract.
2020 Mathematics Subject Classification:  35Q55,35B65,35P25,35B40,47A75,35B30,
81Q10
@@
E-mail: diao@jlu.edu.cn, hadiao@gmail.com
@@ 1. Introduction
1.1. Background and discussion
The Schrödinger equation is one of the fundamental dispersive models in mathematical physics. In its time-dependent form, it describes the evolution of complex wave functions and provides the basic framework for the analysis of non-relativistic quantum dynamics. From the viewpoint of partial differential equations, its dispersive structure gives rise to a rich analytic theory, including decay estimates, Strichartz estimates, scattering theory, and delicate regularity phenomena; see, for instance, [4, 19]. In many physically relevant settings, one must also incorporate nonlinear self-interaction effects, which leads to nonlinear Schrödinger-type equations. A canonical model is the nonlinear Schrödinger equation
$i\,\partial_{t}u+\Delta u+\kappa\,|u|^{p-1}u=0,\qquad(x,t)\in\mathbb{R}^{d}\times\mathbb{R},$ (1.1)
where $\kappa\in\mathbb{R}$ and $p>1$. The focusing and defocusing regimes exhibit markedly different dynamics, and the equation arises naturally in nonlinear optics, Bose–Einstein condensation, and water wave theory; see [18, 4, 19].
The mathematical theory of Schrödinger equations has developed along several major directions, among which well-posedness, scattering, and singularity formation are particularly prominent. In the focusing regime, finite-time blow-up is a central phenomenon. Classical results show that sufficiently localized smooth initial data may generate singular behavior in supercritical settings [10], while in critical regimes one may even prescribe blow-up at finitely many spatial points [15, 8]. On the other hand, localization phenomena also play an important role in the spectral and dynamical theory of Schrödinger operators. A representative example is Anderson localization [2], where random or quasi-periodic structures may force eigenfunctions to concentrate strongly near certain spatial regions; see, e.g., [6, 16, 20, 23]. These two directions, singularity formation and localization, reflect in distinct manners the capacity of Schrödinger dynamics to generate highly nontrivial spatial concentration.
The purpose of the present work is to establish a distinct concentration mechanism for a broad class of linear and nonlinear Schrödinger-type equations. More precisely, we show that, over any prescribed finite time interval $[0,T]$, one can construct smooth initial and/or boundary data such that the corresponding solution remains locally $H^{3}$-regular, yet exhibits extraordinarily large spatial gradients in arbitrarily small neighborhoods of any prescribed finite collection of boundary points. This phenomenon is fundamentally different from classical blow-up: no loss of regularity occurs, and the large-gradient behavior is both controlled and customizable. It is also different in nature from Anderson localization, which is induced by randomness or spectral disorder. Here the localization is deterministic and constructive, produced through carefully designed inputs and realized within globally regular Schrödinger evolutions.
A further feature of our construction is that the region of amplification becomes quantitatively small as the prescribed gradient threshold increases. More precisely, if $\mathcal{M}$ denotes the target size of the gradient, then the measure of the set on which the gradient exceeds $\mathcal{M}$ decays like $O(\mathcal{M}^{-d})$; see Theorems 1.1 and 1.2. Thus, increasingly large gradients are achieved on increasingly small spatial sets. In this sense, our results reveal a quantitative concentration mechanism that is compatible with the dispersive character of Schrödinger dynamics
1.2. Mathematical setup and summary of main results
We focus on the mathematical models for both linear and nonlinear Schrödinger equations in two and three spatial dimensions in this paper, where the mechanism for these mathematical models will be elaborated in more detail later. Consider that $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{d}$, $d=2,3$, and $0<T<\infty$. Furthermore, assume that $D$ is a bounded Lipschitz domain such that $D\Subset\Omega$, and the region $\Omega\setminus\overline{D}$ is connected. Throughout this paper, $\overline{\cdot}$ denotes the complex conjugate and $\top$ denotes the transpose. We define the partial differentiation operator $\mathcal{L}_{1}$ as follows:
$\mathcal{L}_{1}=\nabla\cdot(\mathscr{A}_{1}\nabla),$
where the complex-valued coefficient matrix $\mathscr{A}_{1}=(a_{1}^{ij}(\mathbf{x}))_{i,j=1}^{d}$ satisfies the Hermitian property $\overline{\mathscr{A}}_{1}^{\top}(\mathbf{x})=\mathscr{A}_{1}(\mathbf{x}),~\forall\mathbf{x}\in\Omega$, and belongs to the regularity class $\mathscr{A}_{1}\in W^{2,\infty}(\Omega)\cap C^{1,1}(\overline{\Omega})$. The operator is assumed to be strongly elliptic in the sense that there exists a constant $\theta>0$ such that:
$\overline{\xi}^{\top}\mathscr{A}_{1}(\mathbf{x})\xi\geq\theta|\xi|^{2}$
for a.e. $\mathbf{x}\in\Omega$ and for all $\xi\in\mathbb{C}^{d}$. Furthermore, assume that $\mathscr{A}_{1}-\mathbf{I}$ has compact support in $D$, namely,
$\operatorname{supp}(\mathscr{A}_{1}-\mathbf{I})\subset D,$
@@ 1.2. Mathematical setup and summary of main results
where $\mathbf{I}$ is the $d$-dimensional identity matrix. Under these geometric and structural assumptions, let the unknown function $u:\Omega\times[0,T]\rightarrow\mathbb{C}$ satisfy the following homogeneous initial-boundary value problem for the linear Schrödinger equation:
@@
$\begin{cases}\mathrm{i}\partial_{t}u+\mathcal{L}_{1}u=0&\text{in }\Omega\times(0,T],\\ u(\mathbf{x},0)=\phi(\mathbf{x})&\text{in }\Omega,\\ u(\mathbf{x},t)=\psi(\mathbf{x},t)&\text{on }\partial\Omega\times[0,T],\end{cases}$ (1.2)
where $\mathrm{i}$ denotes $\sqrt{-1}$. Here the initial data $\phi\in C^{\infty}{(\Omega)}$, and boundary data $\psi\in C^{\infty}([0,T];\\ C(\partial\Omega))$ will be specified in Section 3.
Similarly, for the Cauchy problem of a nonlinear Schrödinger-type equation in the full space $\mathbb{R}^{d}$, we consider a bounded Lipschitz domain $D\subset\mathbb{R}^{d}$ with a connected complement. For any given time interval $T\in\mathbb{R}_{+}$, we assume that the zero-order term $c(\mathbf{x},t)$ satisfies:
$c\in L^{\infty}\bigl(0,T;W^{3,\infty}(\mathbb{R}^{d})\bigr),$
and its spatial support is strictly contained in $D$ for all $t\in[0,T]$, i.e.,
$\operatorname{supp}c(\cdot,t)\subset D,\qquad\forall t\in[0,T].$
We introduce the time-dependent partial differentiation operator $\mathcal{L}_{2}$ as follows:
$\mathcal{L}_{2}=\nabla\cdot(\mathscr{A}_{2}\nabla)+c,$
where the complex-valued coefficients $\mathscr{A}_{2}(\mathbf{x},t)=\big(a_{2}^{ij}(\mathbf{x},t)\big)_{i,j=1}^{d}$ are given by the piecewise definition:
$\mathscr{A}_{2}(\mathbf{x},t)=\begin{cases}\mathcal{B}(t),&(\mathbf{x},t)\in\overline{D}\times[0,T],\\[4.0pt] \mathbf{I},&(\mathbf{x},t)\in(\mathbb{R}^{d}\setminus\overline{D})\times[0,T].\end{cases}$
Here $\mathcal{B}(t)=\big(b^{ij}(t)\big)_{i,j=1}^{d}$ is a Hermitian matrix such that $\mathcal{B}\in L^{\infty}[0,T]$. Moreover, the matrix $\mathcal{B}(t)$ satisfies the uniform ellipticity condition:
with $\theta\in\mathbb{R}_{+}$.
@@
Let the unknown function $U:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{C}$ satisfy the following Cauchy problem for a nonlinear Schrödinger equation:
@@
$\begin{cases}\mathrm{i}\partial_{t}U+\mathcal{L}_{2}U+N(U)=0&\text{in }\mathbb{R}^{d}\times(0,T],\\ U(\mathbf{x},0)=\Phi(\mathbf{x})&\text{in }\mathbb{R}^{d},\end{cases}$ (1.3)
where the nonlinearity $N:\mathbb{C}\to\mathbb{C}$ is a complex polynomial defined as:
$N(U)=\sum_{k=2}^{l_{0}}\alpha_{k}U^{k},$ (1.4)
with $l_{0}\geq 2$ being a fixed integer. The complex-valued coefficients $\alpha_{k}$ satisfy the smoothness and localization conditions:
with their spatial support strictly contained in $D$:
The smooth initial data $\Phi\in C^{\infty}(\mathbb{R}^{d})$ will be specified in Section 3.
We are now in a position to state our main theorems. For the initial-boundary value problem of the linear Schrödinger equation (1.2), we have:
Theorem 1.1.
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial and boundary inputs (as specified in (3.21)), there exists a small parameter $r_{0}$ given by (4.14) such that problem (1.2) admits a unique solution $u$ satisfying
$u\in L^{\infty}\big([0,T];H^{3}_{\mathrm{loc}}(\Omega)\big),$
with the following properties:
@@
    Interior improved Hölder regularity: for any bounded Lipschitz domain $\Omega^{\prime}\Subset\Omega$,
@@
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla u(\mathbf{x},t)|>\mathcal{M}\right\}=O(\mathcal{M}^{-d})\qquad\text{as }\mathcal{M}\to\infty,d=2,3.$
Similarly, for the Cauchy problem of the nonlinear Schrödinger-type equation (1.3), we have:
@@
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial inputs (as specified in (3.22)), there exists a small parameter $r_{0}$ given by (5.23) such that problem (1.3) admits a unique solution $U$ satisfying
@@
with the following properties:
@@
    Hölder regularity: for any bounded Lipschitz domain $\Omega^{\prime}\subset\mathbb{R}^{d}$,
@@
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla U(\mathbf{x},t)|>\mathcal{M}\right\}=O(\mathcal{M}^{-d})\qquad\text{as }\mathcal{M}\to\infty,d=2,3.$
@@
In Theorems 1.1 and 1.2, we identify and construct a mechanism that produces highly localized gradient amplification of solutions to both the linear and nonlinear Schrödinger-type equations in two and three spatial dimensions. The phenomenon is described by the following two features:
  1. (i)
    For any prescribed set of $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, where $D$ is the support of the anisotropic coefficients $\mathscr{A}_{1}$ in (1.2) or $\mathscr{A}_{2}$ in (1.3), the zero-order term $c$ in (1.3), and the nonlinear term $N$ in (1.3), we rigorously prove—by selecting suitably tuned smooth initial–boundary data for (1.2) or smooth initial data for (1.3)—that there exist localized gradient amplification regions in neighborhoods of each $\mathbf{x}_{i}$ located outside $D$ ($i=1,\ldots,n$). In these regions, the $C^{0}$-norms of the spatial gradients of the solutions, namely $\nabla u(t)$ to (1.2) and $\nabla U(t)$ to (1.3), exceed any prescribed large constant $\mathcal{M}$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\alpha}$-norm of the solution in these regions and the $C^{1,\alpha}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$. These properties are established in Theorems 1.1 and 1.2, where the time horizon $T\in\mathbb{R}_{+}$ is arbitrary.
  2. (ii)
To the best of our knowledge, the present work constitutes the first systematic and customized construction of solutions that exhibit localized gradient amplification for Schrödinger-type equations across both linear and nonlinear regimes.
@@
The mechanism developed in this work for producing customized solutions with localized gradient amplification exhibits several fundamental features:
  • (i)
    The construction is rooted in spectral analysis, specifically leveraging the transmission eigenfunctions associated with the interior transmission problem [5] and their approximation via Herglotz wave functions. By exploiting these spectral tools, we construct a precisely tuned auxiliary function $u_{0}$ (see Theorem 3.1), which serves as the essential building block for prescribing the initial and boundary inputs in both the linear and nonlinear regimes. Detailed construction procedures are provided in Section 3.
  • (ii)
@@
    A crucial aspect of our results is that the gradient amplification occurs within the framework of classical derivatives. Both the solution $u(t)$ to the linear problem (1.2) and $U(t)$ to the nonlinear problem (1.3) exhibit local $H^{3}$-regularity for almost every $t\in[0,T]$. By the Sobolev embedding theorem, this ensures that the solutions possess $C^{1,\frac{1}{2}}$-regularity in the neighborhoods of the amplification points. Consequently, the spatial gradients remain well-defined in the classical sense even as their magnitudes exceed the arbitrarily large threshold $\mathcal{M}$.
  • (iii)
@@
    The localized gradient amplification is not a transient phenomenon but persists over any prescribed time interval $[0,T]$ for $T\in\mathbb{R}_{+}$. While this follows from classical well-posedness in the linear case, the nonlinear setting requires a more refined analysis. By carefully tuning the initial data and applying a Banach fixed-point argument, we ensure the stability and existence of the solution $U(t)$ over an arbitrarily large time horizon $T$, as further detailed in Remark 5.3 and the proof of Theorem 1.2.